OpenStudy (anonymous):

Input 1. F(3) Input 2. F(4) F (x) = 6(x) + 2(x) Can anyone help me answer this?

OpenStudy (anonymous):

Input 1 F(3) = 6*3 + 2*3 + 18 + 6 = 24 Input 2 F(4) = 6*4 + 2*4 = 32

OpenStudy (anonymous):

Whoops i meant 6*3 + 2*3 = 18 + 6 = 24

OpenStudy (anonymous):

oh ok, that looks better lol, now it makes sense thank you!